OpenStudy (anonymous):

Use a both elimination and substitution to solve this system of linear equations. Type your answer as an ordered triplet, like this: (-2, 4, 13) 2x−10y+1=z 4x + y = z + 18 -2x - 8z = y - 1

OpenStudy (anonymous):

Ok, lets start by solving one of the equations for one of the variables. Which variable would you like to start with?

OpenStudy (anonymous):

the first set

OpenStudy (anonymous):

If I said a good place to start would be the variable z, why do you think that would be a good idea?

OpenStudy (anonymous):

cause its easiest?

OpenStudy (anonymous):

Right! It's already solved for. 2x−10y+1 = z z = 2x−10y+1 Ok so we know what z is equal too. Let's use this information to plug substitute what we got for z into one of the other equations.

OpenStudy (anonymous):

The second equation has z by itself in there. Let's plug in 2x−10y+1 for z since z = 2x−10y+1 We would get from 4x + y = z + 18 To 4x + y = (2x - 10y +1) +18 Right? Do you have any questions so far?

OpenStudy (anonymous):

not so far

OpenStudy (anonymous):

Actually I'm making this harder than it needs to be. Let's actually solve for y in the second equation since it is by itself. 4x + y = z + 18 Comes to y = -4x + z + 18 Right? So we have, z = 2x−10y+1 y = -4x + z + 18

OpenStudy (anonymous):

The order of steps can really make things complicated or easy depending on which route you take. The way I was doing it would have worked but would have been a nightmare to work through haha.

OpenStudy (anonymous):

Are you with me so far? When you are, we can continue.